OpenStudy (anonymous):

Use chunking and u-sustitution to find the zeros of f(x)=(x^4)+5(x^2)-36.

sam (.sam.):

(x^4)+5(x^2)-36 (x^2)^2+5(x^2)-36 u=x^2 u^2+5u-36 (u-4)(u+3)=0 u=4 u=-3 ------------ x^2=4 x=2,-2 ------------- x^2=-3 np